Course Announcement – Spring 1998

Math 242 – Elementary Differential Equations Professors Miller (Section 1) and Meade (Section 3)

\underline{Section 1} (schedule code: 204838)  \underline{Section 3} (schedule code: 204878)
Professor  Matt Miller  Doug Meade
Office  LeConte 300I  LeConte 300E
Phone  777-3690  777-6183

 e-mail  miller@math.sc.edu               meade@math.sc.edu
 WWW URL http://www.math.sc.edu/˜ miller/  http://www.math.sc.edu/˜ meade/

Meeting Time MWF 9:05 – 9:55  TTh 11:00 – 12:15
Text  Differential Equations, Blanchard, Devaney, & Hall, Brooks/Cole, 1998

Course Description

Sections 1 and 3 of Math 242 will be taught using the new Blanchard, Devaney, and Hall text from Brooks/Cole. The selection of this textbook was based largely on its unique and exciting approach to differential equations and their applications. For example, instead of concentrating almost exclusively on methods that produce closed-form solutions, the primary emphasis is on qualitative, geometric, and numerical techniques for the analysis of differential equations. This approach facilitates an in-depth discussion of systems and nonlinear models with an emphasis on phase-plane analysis and questions concerning stability.

Specific skills that will be developed in this course include an understanding of the geometry of solutions to differential equations, the ability to work with multiple representations of solutions, techniques for the construction and analysis of models of physically realistic situations, and the ability to communicate qualitative behavior of solutions in writing.

This course will be taught in the new student computer facility in the Department of Mathematics (LC 303A, a networked PC lab). Students will be introduced to, and expected to learn to use, Maple. Maple is a computer algebra system that is particularly well-suited to the approach used with this text, and is available in all of the science and engineering computer labs on campus.

Additional information (including this document) can be found on the WWW. To access this information, point your browser to the URL

                 http://www.math.sc.edu/˜ meade/math242-S98/